# references.bib

### v3

@@ 6.3. $\mathrm{GL}_{2}(\mathbf{C})\curvearrowright(\mathbf{C}^{2}\oplus(\mathbf{C}^{2})^{\vee})^{g}$
We generalize the previous situation by considering the diagonal action of $\mathrm{GL}_{2}(\mathbf{C})$ on $(\mathbf{C}^{2}\oplus(\mathbf{C}^{2})^{\vee})^{g}$. We can describe the induction maps as follows, analogously to §6.2.
@@
We have $\mathcal{P}_{\lambda_{0}}=\bigoplus_{j=0}^{g-2}\mathbf{Q}(x_{1}+x_{2})^{j}$, $\mathcal{P}_{\lambda_{1}}=\bigoplus_{j=0}^{g-1}\mathbf{Q}x_{2}^{j}$, $\mathcal{P}_{\lambda_{2}}=\mathbf{Q}$ and the cohomological integrality isomorphism reads
@@ 6.4. $\mathrm{SL}_{2}(\mathbf{C})\curvearrowright\mathbf{C}^{2}$
$\begin{matrix}\lambda_{1}&\colon&\mathbf{G}_{\mathrm{m}}&\rightarrow&T\\ &&v&\mapsto&(v,v^{-1}).\end{matrix}$
@@
We have $\mathcal{P}_{\lambda_{1}}=\mathbf{Q}$ and $\mathcal{P}_{\lambda_{0}}=\{0\}$ and the cohomological integrality isomorphism is
$\begin{matrix}\mathbf{Q}[x^{2}]&\rightarrow&\mathbf{Q}[x]\\ f&\mapsto&f\,.\end{matrix}$
@@ 6.5. $\mathrm{SL}_{2}(\mathbf{C})\curvearrowright\mathbf{C}^{d}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{1},\lambda_{0}}&\colon&\mathbf{Q}[x]&\rightarrow&\mathbf{Q}[x^{2}]\\ &&f&\mapsto&C_{d}x^{\lfloor\frac{d-1}{2}\rfloor}f(x)+C_{d}(-x)^{\lfloor\frac{d-1}{2}\rfloor}f(-x).\end{matrix}$
We have $\mathcal{P}_{\lambda_{1}}=\mathbf{Q}$ and $\mathcal{P}_{\lambda_{0}}=\mathbf{Q}[x^{2}]_{\deg<\lfloor\frac{d-1}{2}\rfloor}$ where $C_{d}\in\mathbf{Q}$ is some constant depending only on $d$.
@@
$\begin{matrix}\mathcal{P}_{\lambda_{0}}\oplus\mathbf{Q}[x]^{\varepsilon_{d}}&\rightarrow&\mathbf{Q}[x]\\ f,g&\mapsto&f+2C_{d}x^{\lfloor\frac{d-1}{2}\rfloor}g\end{matrix}$
where $\mathbf{Q}[x]^{\varepsilon_{d}}=\mathbf{Q}[x^{2}]$ if $\lfloor\frac{d-1}{2}\rfloor$ is even, and $\mathbf{Q}[x]^{\varepsilon_{d}}=x\mathbf{Q}[x^{2}]$ if $\lfloor\frac{d-1}{2}\rfloor$ is odd.
@@
We note, for example in §6.2 and also §6.4, that the integrality morphisms are not isomorphisms when we consider cohomology with integral coefficients instead of rational coefficients. In §6.2, this comes from the presence of the factor $2$ in the cohomological integrality morphism.
6.6. $\mathrm{SL}_{2}(\mathbf{C})\curvearrowright(\mathfrak{sl}_{2}(\mathbf{C}))^{g}$
Let $g\geqslant 0$. We consider the diagonal adjoint action of $\mathrm{SL}_{2}(\mathbf{C})$ on $V\coloneqq(\mathfrak{sl}_{2}(\mathbf{C}))^{g}$. Again, $\mathscr{P}_{V}=\{\overline{\lambda_{1}},\overline{\lambda_{1}}\}$ with $\lambda_{0},\lambda_{1}$ as in §6.4. The induction map reads
@@ 6.6. $\mathrm{SL}_{2}(\mathbf{C})\curvearrowright(\mathfrak{sl}_{2}(\mathbf{C}))^{g}$
We have $V^{\lambda_{1}}\cong\mathbf{C}^{g}$ with the trivial action of $T$. Therefore, $\mathcal{P}_{\lambda_{1}}=\mathbf{Q}$ and $\mathcal{P}_{\lambda_{0}}=\mathbf{Q}[x^{2}]_{\deg<g-1}$. The cohomological integrality isomorphism is
$\begin{matrix}\mathbf{Q}[x]^{\epsilon_{g}}\oplus\mathcal{P}_{0}&\rightarrow&\mathbf{Q}[x^{2}]\ f,g&\mapsto&g(x)+2^{g}x^{g-1}f(x)\end{matrix}$
where $\mathbf{Q}[x]^{\epsilon_{g}}=\mathbf{Q}[x^{2}]$ if $g-1$ is even and $\mathbf{Q}[x]^{\epsilon_{g}}=x\mathbf{Q}[x^{2}]$ is $g-1$ is odd.
@@
This example is an example for which $G_{\lambda_{0}}=\{\pm 1\}$ is a finite group. We see that it does not affect the cohomological integrality isomorphism.
@@ 6.7. $G\curvearrowright\mathfrak{g}$
Let $G$ be a reductive group and $\mathfrak{g}$ its Lie algebra with the adjoint action. Then, $\mathscr{P}_{\mathfrak{g}}$ is inn one-to-one correspondence with the set of standard Levi subgroups of $G$ and $\mathscr{P}_{\mathfrak{g}}/W$ is in bijection with conjugacy classes of Levi subgroups of $\mathfrak{g}$. In this case, $\mathbf{Q}[\mathfrak{t}]^{W}$ is a polynomial algebra and for $\lambda\preceq\mu\in X_{*}(T)$, the induction map is
@@ 6.8. Topology of the algebra of invariants
The sheafified version of our cohomological integrality theorem, which is the subject of a forthcoming work, would give an algorithm to compute the intersection cohomology of the GIT quotients $V_{d}/\!\!/\mathrm{SL}_{2}$ where $V_{d}$ is the $d$-dimensional irreducible representation of $\mathrm{SL}_{2}(\mathbf{C})$.
For $d\in\mathbf{N}$, one can define an $\mathrm{SL}_{2}(\mathbf{C})$-action on the space of binary forms of the form
$b(x,y)=\sum_{j=0}^{d}a_{i}x^{i}y^{d-i}$
by making $\mathrm{SL}_{2}(\mathbf{C})$ act on the coordinates $(x,y)$ via the natural representation. This induces an action of $\mathrm{SL}_{2}(\mathbf{C})$ on the polynomial algebra $\mathbf{C}[a_{i},1\leqslant i\leqslant d]$. One wants to compute the invariant ring $\mathbf{C}[a_{i},1\leqslant i\leqslant d]^{\mathrm{SL}_{2}(\mathbf{C})}$. Let $V_{d}$ be the irreducible $d$-dimensional representation of $\mathrm{SL}_{2}(\mathbf{C})$. In modern terms, we try to compute the geometric invariant theory quotient $V_{d}/\!\!/\mathrm{SL}_{2}(\mathbf{C})$. We may give answers to this problem in small degrees.
@@ 6.8.1. Invariants of linear binary forms
Let $V_{2}=\mathbf{C}^{2}$ be the natural representation of $\mathrm{SL}_{2}(\mathbf{C})$. Then, we calculate easily $V_{2}/\!\!/\mathrm{SL}_{2}(\mathbf{C})=\mathrm{pt}$ since the action of $\mathrm{SL}_{2}(\mathbf{C})$ on $V_{2}$ has two orbits, $\{0\}$ and $V_{2}\smallsetminus\{0\}$, of those only $\{0\}$ is closed. The stabilizer of the closed orbit $\{0\}$ is $\mathrm{SL}_{2}(\mathbf{C})$, which is not finite. Therefore, the stable locus is empty.
@@ 6.8.2. Invariants of binary quadratic forms
$D=b^{2}-4ac.$
Therefore, we have $V_{2}/\!\!/\mathrm{SL}_{2}(\mathbf{C})\cong\mathbf{C}$. Since $\dim V_{3}/\mathrm{SL}_{2}(\mathbf{C})=0$ and $\dim\mathbf{C}=1$, the stable locus is empty.
@@ 6.8.3. Invariants of binary cubic forms
$D=3b^{2}c^{2}+6abcd-4b^{3}d-4c^{3}a-a^{2}d^{2}.$
Therefore, $V_{4}/\!\!/\mathrm{SL}_{2}(\mathbf{C})\cong\mathbf{C}$. Moreover, since $\dim V_{4}/\mathrm{SL}_{2}(\mathbf{C})=1=\dim\mathbf{C}$, the stable locus is non-empty. By $\mathbf{C}^{*}$-equivariance, the stable locus is the open locus of binary cubic forms of non-zero discriminant.
We notice that $\dim\mathrm{IH}(\mathbf{C}^{2})=1=\dim\mathcal{P}_{\lambda_{0}}=1$ where $\mathcal{P}_{\lambda_{0}}$ is given in §6.5 for $d=4$.
@@ 6.8.4. Invariants of binary quartic forms
Let $V_{5}$ be the irreducible $5$-dimensional representation of $\mathrm{SL}_{2}(\mathbf{C})$. The ring of invariants of binary quartic forms has two algebraically independent generators $i,j$. Therefore, $V_{5}/\!\!/\mathrm{SL}_{2}(\mathbf{C})\cong\mathbf{C}^{2}$. Moreover, $\dim V_{5}/\mathrm{SL}_{2}(\mathbf{C})=2=\dim\mathbf{C}^{2}$ and therefore, the stable locus is non-empty.
6.8.5. Invariants of binary quintic forms
Let $V_{6}$ be the irreducible $6$-dimensional representation of $\mathrm{SL}_{2}(\mathbf{C})$. There are four polynomial invariants $I_{4},I_{8},I_{12},I_{18}$, of respective degrees $4,8,12,18$. The first three $I_{4},I_{8},I_{12}$ are algebraically independent while the square of the last can be expressed in terms of them:
$I_{18}^{2}=-144I_{12}^{3}+(I_{4}^{3}-72I_{4}I_{8})I_{12}^{2}+(24I_{8}^{2}-6I_{4}^{2}I_{8}^{2})I_{12}+9I_{4}I_{8}^{4}.$
This realizes the GIT quotient $V_{6}/\!\!/\mathrm{SL}_{2}(\mathbf{C})$ as a double cover of $\mathbf{C}^{3}$ ramified over the hypersurface $-144I_{12}^{3}+(I_{4}^{3}-72I_{4}I_{8})I_{12}^{2}+(24I_{8}^{2}-6I_{4}^{2}I_{8}^{2})I_{12}+9I_{4}I_{8}^{4}=0$.
6.8.6. Invariant of binary sextic forms
Let $V_{7}$ be the $7$-dimensional irreducible representation of $\mathrm{SL}_{2}(\mathbf{C})$. The ring of polynomial invariants of binary sextic forms has $5$ generators $I_{2},I_{4},I_{6},I_{10},I_{15}$ of respective degrees $2,4,6,10,15$. The generators $I_{2},I_{4},I_{6},I_{10}$ are algebraically independent and $I_{1}^{2}$ can be expressed in terms of $I_{2},I_{4},I_{6},I_{10}$. This realizes $V_{7}/\!\!/\mathrm{SL}_{2}(\mathbf{C})$ has a double cover of $\mathbf{C}^{4}$ ramified over a hypersurface. Moreover, since $\dim V_{7}/\!\!/\mathrm{SL}_{2}(\mathbf{C})=4=\dim\mathbf{C}^{4}$, the stable locus is non-empty.
[…]

### v6

@@ 6.3. $\mathrm{GL}_{2}(\mathbf{C})\curvearrowright(\mathbf{C}^{2}\oplus(\mathbf{C}^{2})^{\vee})^{g}$
We generalize the previous situation by considering the diagonal action of $\mathrm{GL}_{2}(\mathbf{C})$ on $(\mathbf{C}^{2}\oplus(\mathbf{C}^{2})^{\vee})^{g}$ ($g\geqslant 0$). The set $\mathscr{P}_{V}/W$ is the same as in §6.2. The fixed point sets $V^{\lambda_{0}},V^{\lambda_{1}}$ and $V^{\lambda_{2}}$ can be computed similarly and the Levi subgroups $G^{\lambda_{0}},G^{\lambda_{1}}$ and $G^{\lambda_{2}}$ are the ones given in §6.2. We can describe the induction maps as follows, analogously to §6.2.
@@
$\begin{matrix}\mathrm{Ind}_{\lambda_{2},\lambda_{3}}&\colon&\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1}+x_{2},x_{1}x_{2}]\\ &&f&\mapsto&\frac{f(x_{1},x_{2})-f(x_{2},x_{1})}{x_{1}-x_{2}}\end{matrix}$
We have $\mathcal{P}_{\lambda_{0}}=\bigoplus_{j=0}^{g-2}\mathbf{Q}(x_{1}+x_{2})^{j}$, $\mathcal{P}_{\lambda_{1}}=\bigoplus_{j=0}^{g-1}\mathbf{Q}x_{2}^{j}$, $\mathcal{P}_{\lambda_{2}}=\mathbf{Q}$ and $\mathcal{P}_{\lambda_{3}}=0$ and the cohomological integrality isomorphism reads
@@ 6.4. $\mathrm{SL}_{2}(\mathbf{C})\curvearrowright\mathbf{C}^{2}$
$\begin{matrix}\lambda_{1}&\colon&\mathbf{G}_{\mathrm{m}}&\rightarrow&T\\ &&v&\mapsto&(v^{-1},v)\,.\end{matrix}$
@@
The factor $\frac{1}{2}$ comes from the fact that the $T$-weight of $\mathfrak{g}^{\lambda_{1}>0}=\begin{pmatrix}0&*\\ 0&0\end{pmatrix}$ is $2$. We have $\mathcal{P}_{\lambda_{1}}=\mathbf{Q}$ and $\mathcal{P}_{\lambda_{0}}=\{0\}$ and the cohomological integrality isomorphism is
$\begin{matrix}\mathbf{Q}[x^{2}]&\rightarrow&\mathbf{Q}[x^{2}]\\ f&\mapsto&f\,.\end{matrix}$
@@ 6.5. $\mathrm{SL}_{2}(\mathbf{C})\curvearrowright\mathbf{C}^{d}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{1},\lambda_{0}}&\colon&\mathbf{Q}[x]&\rightarrow&\mathbf{Q}[x^{2}]\\ &&f&\mapsto&C_{d}x^{\lfloor\frac{d}{2}\rfloor-1}f(x)+C_{d}(-x)^{\lfloor\frac{d}{2}\rfloor-1}f(-x)\,.\end{matrix}$
We have $\mathcal{P}_{\lambda_{1}}=\mathbf{Q}$ and $\mathcal{P}_{\lambda_{0}}=\mathbf{Q}[x^{2}]_{\deg<2(\lfloor\frac{d}{2}\rfloor-1)}$ where $C_{d}\in\mathbf{Q}$ is some constant depending only on $d$ (note that $\deg(x)=2$ in the cohomological grading).
@@
$\begin{matrix}\mathcal{P}_{\lambda_{0}}\oplus\mathbf{Q}[x]^{\varepsilon_{d}}&\rightarrow&\mathbf{Q}[x^{2}]\\ (f,g)&\mapsto&f+2C_{d}x^{\lfloor\frac{d}{2}\rfloor-1}g\end{matrix}$
where $\mathbf{Q}[x]^{\varepsilon_{d}}=\mathbf{Q}[x^{2}]$ if $\lfloor\frac{d}{2}\rfloor-1$ is even, and $\mathbf{Q}[x]^{\varepsilon_{d}}=x\mathbf{Q}[x^{2}]$ if $\lfloor\frac{d}{2}\rfloor-1$ is odd.
@@
We note, for example in §6.2 and also §6.4, that the integrality morphisms are not isomorphisms when we consider cohomology with integral coefficients instead of rational coefficients. In §6.2, this comes from the presence of the factor 22 in the cohomological integrality morphism.
$6.$ $\mathrm{SL}_{2}(\mathbf{C})\curvearrowright(\mathfrak{sl}_{2}(\mathbf{C}))^{g}$
Let $g\geqslant 0$. We consider the diagonal adjoint action of $\mathrm{SL}_{2}(\mathbf{C})$ on $V\coloneqq(\mathfrak{sl}_{2}(\mathbf{C}))^{g}$. Again, $\mathscr{P}_{V}=\{\overline{\lambda_{0}},\overline{\lambda_{1}}\}$ with $\lambda_{0},\lambda_{1}$ as in §6.4. The induction map reads
@@
We have $V^{\lambda_{1}}\cong\mathbf{C}^{g}$ with the trivial action of $T$. Therefore, $\mathcal{P}_{\lambda_{1}}=\mathbf{Q}$ and $\mathcal{P}_{\lambda_{0}}=\mathbf{Q}[x^{2}]_{\deg<2(g-1)}$. The cohomological integrality isomorphism is
$\begin{matrix}\mathbf{Q}[x]^{\varepsilon_{g}}\oplus\mathcal{P}_{0}&\rightarrow&\mathbf{Q}[x^{2}]\\ (f,g)&\mapsto&g(x)+2^{g}x^{g-1}f(x)\end{matrix}$
where $\mathbf{Q}[x]^{\varepsilon_{g}}=\mathbf{Q}[x^{2}]$ if g1g-1 is even and $\mathbf{Q}[x]^{\varepsilon_{g}}=x\mathbf{Q}[x^{2}]$ is g1g-1 is odd.
@@
In this example, $G_{\lambda_{0}}=\{\pm 1\}$ is a finite group. We see that it does not affect the cohomological integrality isomorphism.
6.7. $G\curvearrowright\{0\}$
In this section, we present the cohomological integrality isomorphism for the trivial representation of a reductive group $G$. Then, $\mathscr{P}_{\{0\}}/W$ is in bijection with the set of conjugacy classes of Levi subgroups of $G$. If $\lambda\in X_{*}(T)$ is a general cocharacter, then the induction morphism
$\mathrm{Ind}_{\lambda}\colon\mathrm{H}^{*}(\mathrm{pt}/T)\rightarrow\mathrm{H}^{*}(\mathrm{pt}/G)$
is surjective. Indeed, we have
$\mathrm{Ind}_{\lambda}(f)=\sum_{w\in W}w\left(\frac{f}{\prod_{\begin{subarray}{c}\alpha\in\mathcal{W}(\mathfrak{g})\\ \langle\lambda,\alpha\rangle<0\end{subarray}}\alpha}\right).$
Therefore, $f=\mathrm{Ind}_{\lambda}\left(\frac{1}{\#W}f\prod_{\begin{subarray}{c}\alpha\in\mathcal{W}(\mathfrak{g})\\ \langle\lambda,\alpha\rangle<0\end{subarray}}\alpha\right)$.
This implies that $\mathcal{P}_{\lambda}=\{0\}$ if $G^{\lambda}$ is not the maximal torus of $G$ and $\mathcal{P}_{\lambda}=\mathbf{Q}$ if $G^{\lambda}=T$. In particular, the cohomological integrality reads (for a choice of generic $\lambda\in X_{*}(T)$):
[…]
where $\varepsilon_{W}(w)=(-1)^{\ell(w)}$ is the sign character of the Coxeter group $W$; $\ell(w)$ denotes the length of $w\in W$. This morphism is indeed an isomorphism since $\prod_{\begin{subarray}{c}\alpha\in\mathcal{W}(\mathfrak{g})\\ \langle\lambda,\alpha\rangle<0\end{subarray}}\alpha$ is a generator of the $\mathbf{Q}[\mathfrak{t}]^{W}$-module of anti-invariant polynomials.
@@ 6.8. $G\curvearrowright\mathfrak{g}$
Let $G$ be a reductive group and $\mathfrak{g}$ its Lie algebra with the adjoint action. Then, $\mathscr{P}_{\mathfrak{g}}$ is in one-to-one correspondence with the set of standard Levi subgroups of $G$ and $\mathscr{P}_{\mathfrak{g}}/W$ is in bijection with conjugacy classes of Levi subgroups of $\mathfrak{g}$. In this case, $\mathbf{Q}[\mathfrak{t}]^{W}$ is a polynomial algebra and for $\lambda\preceq\mu\in X_{*}(T)$, the induction map is
@@ 6.9. Topology of the algebra of invariants
The strong version of the sheafified upgrade of our cohomological integrality theorem, which is the subject of a forthcoming work, would give an algorithm to compute the intersection cohomology of the GIT quotients V//GV/\!\!/G for symmetric representations $V$ of a reductive group $G$.
Conjecture 6.3 (Strong integrality conjecture).
Let $\lambda\in X_{*}(T)$. Then, there is a canonical identification $\mathcal{P}_{\lambda}\cong\mathrm{IH}^{*}(V^{\lambda}/\!\!/G^{\lambda})$ if a general closed orbit of $G^{\lambda}/G_{\lambda}$ inside $V^{\lambda}$ has finite stabilizer in $G^{\lambda}/G_{\lambda}$, and $\mathcal{P}_{\lambda}=0$ otherwise.
In particular, for $G=\mathrm{SL}_{2}(\mathbf{C})$, it gives a computation of the intersection cohomology of $V_{d}/\!\!/\mathrm{SL}_{2}$ where $V_{d}$ is the $d$-dimensional irreducible representation of $\mathrm{SL}_{2}(\mathbf{C})$. In this section, we verify this strong integrality conjecture for some representations of $\mathrm{SL}_{2}(\mathbf{C})$. More precisely, we check that for small values of $d$, when a general closed orbit in $V_{d}$ has finite stabilizers, then $\dim\mathrm{IH}(V_{d}/\!\!/\mathrm{SL}_{2}(\mathbf{C}))=\dim\mathcal{P}_{0}$, where $\mathcal{P}_{0}$ is the vector space given by Theorem 1.1 for the trivial cocharacter $\lambda=0=\lambda_{0}$.
@@
Remark 6.4.
While this paper was under refereeing process, the work [Bu+25] appears. The authors solve Conjecture 6.3 under an orthogonality assumption for the representation $V$ of $G$: they assume that there is a nondegenerate symmetric $G$-invariant bilinear on $V$. This condition is stronger than symmetricity (and a fortiori, weak symmetricity) we assume in the present paper, as demonstrated for example by the natural representation $\mathbf{C}^{2n}$ of the symplectic group $\mathrm{Sp}(2n)$. This example can be treated by a direct approach. In combination with [Kno07], it is possible to prove Conjecture 6.3 for any symmetric representation $V$ of $G$ such that $\mathbf{C}[V]^{G}=\mathbf{C}$, that is $V/\!\!/G=\mathrm{pt}$.
@@ 6.9.1. Invariants of linear binary forms
Let $V_{2}=\mathbf{C}^{2}$ be the natural representation of $\mathrm{SL}_{2}(\mathbf{C})$. Then, we calculate easily $V_{2}/\!\!/\mathrm{SL}_{2}(\mathbf{C})=\mathrm{pt}$ since the action of $\mathrm{SL}_{2}(\mathbf{C})$ on $V_{2}$ has two orbits, $\{0\}$ and $V_{2}\smallsetminus\{0\}$, of those only $\{0\}$ is closed. The stabilizer of the closed orbit $\{0\}$ is $\mathrm{SL}_{2}(\mathbf{C})$, which is not finite. Therefore, the stable locus is empty. One expects that $\mathcal{P}_{\lambda_{0}}=0$. This is confirmed by §6.4.
@@ 6.9.2. Invariants of binary quadratic forms
$D=b^{2}-4ac\,$,
see [Sch13, Satz 1.9, 2.8]. Therefore, we have $V_{2}/\!\!/\mathrm{SL}_{2}(\mathbf{C})\cong\mathbf{C}$. Since $\dim V_{3}/\mathrm{SL}_{2}(\mathbf{C})=0$ and $\dim\mathbf{C}=1$, the stable locus is empty. Therefore, we expect that $\mathcal{P}_{\lambda_{0}}=0$, and this is confirmed by §6.5.
@@ 6.9.3. Invariants of binary cubic forms
$D=3b^{2}c^{2}+6abcd-4b^{3}d-4c^{3}a-a^{2}d^{2}\,$,
see [Sch13, Satz 2.8]. Therefore, $V_{4}/\!\!/\mathrm{SL}_{2}(\mathbf{C})\cong\mathbf{C}$. Moreover, since $\dim V_{4}/\mathrm{SL}_{2}(\mathbf{C})=1=\dim\mathbf{C}$, the stable locus is non-empty. By $\mathbf{C}^{*}$-equivariance, the stable locus is the open locus of binary cubic forms of non-zero discriminant.
We notice that $\dim\mathrm{IH}(\mathbf{C})=1=\dim\mathcal{P}_{\lambda_{0}}=1$ where $\mathcal{P}_{\lambda_{0}}$ is given in §6.5 for $d=4$.
@@ 6.9.4. Invariants of binary quartic forms
Let $V_{5}$ be the irreducible $5$-dimensional representation of $\mathrm{SL}_{2}(\mathbf{C})$. The ring of invariants of binary quartic forms has two algebraically independent generators $i,j$ [Sch13, Satz 2.9]. Therefore, $V_{5}/\!\!/\mathrm{SL}_{2}(\mathbf{C})\cong\mathbf{C}^{2}$. Moreover, $\dim V_{5}/\mathrm{SL}_{2}(\mathbf{C})=2=\dim\mathbf{C}^{2}$ and therefore, the stable locus is non-empty. In this case, we expect that $P_{\lambda_{0}}$ is of dimension $1=\dim\mathrm{IH}^{2}(\mathbf{C}^{2})$, and this is again confirmed by §6.5 for $d=5$.
Competing interests
The author has no competing interest to declare.